find the root between (2,3) of x^3-2x-5=0, by using false position method


1
Expert's answer
2022-03-09T14:25:54-0500

According to the false position method

"x_n = x_1 - \frac{x_{n-1}-x_1}{f(x_{n-1}-f(x_1)}f(x_1)"

In our case "f(x) = x^3-2 x-5" , and x1 = 2, x2 = 3

"\\begin{matrix}\nn\t& x_n\t&f(x_n) \\\\\n1&\t2&\t-1 \\\\\n2&\t3&\t16 \\\\\n3&\t2.058823529&\t-0.390799919 \\\\\n4&\t2.096558637&\t0.022428062 \\\\\n5&\t2.094440519&\t-0.001238424 \\\\\n6&\t2.094557621&\t6.85303\\cdot10{-5} \\\\\n7&\t2.094551142&\t-3.79179\\cdot10^{-6} \\\\\n8&\t2.0945515&\t2.09802\\cdot 10^{-7} \\\\\n\\end{matrix}"